# A High Efficient and Scalable Obstacle-Avoiding VLSI Global Routing Flow

### v2

@@ Abstract
Routing is a crucial step in the VLSI design flow. With the advancement of manufacturing technologies, more constraints have emerged in design rules, particularly regarding obstacles during routing, leading to increased routing complexity. Unfortunately, many global routers struggle to efficiently generate obstacle-free solutions due to the lack of scalable obstacle-avoiding tree generation methods and the capability of handling modern designs with complex obstacles and nets. In this work, we propose an efficient obstacle-aware global routing flow for VLSI designs with obstacles. The flow includes a rule-based obstacle-avoiding rectilinear Steiner minimal tree (OARSMT) algorithm during the tree generation phase. This algorithm is both scalable and fast to provide tree topologies avoiding obstacles in the early stage globally. With its guidance, OARSMT-guided and obstacle-aware sparse maze routing are proposed in the later stages to minimize obstacle violations further and reduce overflow costs. Compared to advanced methods on the benchmark with obstacles, our approach successfully eliminates obstacle violations, and reduces wirelength and overflow cost, while sacrificing only a limited number of via counts and runtime overhead.
@@ 1 Introduction
As the number of cells and nets in modern designs has increased dramatically in advanced technology, more complicated obstacles have emerged in global routing to accommodate more complex functionality requirements or to maintain consistency with earlier design phases. These obstacles may arise from various sources, including macro cells, intellectual property (IP) blocks, 3D packaging, pre-routed nets, power networks, etc. However, many global routers struggle to efficiently handle obstacle information and generate obstacle-free solutions at the global routing phase. High-efficiency obstacle-avoiding global routing has become more challenging and more necessary than ever before.
Global routing typically involves several steps, including tree generation, initial routing, and rip-up and reroute phases. Traditional routers that lack specialized methods for avoiding obstacles typically generate rectilinear Steiner minimal trees (RSMTs) for nets, which can be quickly obtained using established algorithms like FLUTE [1]. Without full consideration of obstacles, numerous design-rule violations can occur in the solutions. Although these violations can be addressed in later stages, with more time-consuming algorithms like maze routing, the overall efficiency and effectiveness of the routing process can become limited. To address the obstacle-avoiding RSMT (OARSMT) problem, recent research has introduced several algorithms for single-net solutions, such as complex graph algorithms, machine learning, etc. With more complex obstacle distributions, the complexity can significantly increase, especially for high fan-out nets. Additionally, the number of nets has also risen dramatically in modern designs. These result in scalability problems for most OARSMT algorithms in practicality when applied to large-scale cases in global routing. As a result, most of the previous researches primarily concentrate on the tree generation phase and lack a comprehensive obstacle-avoiding global routing flow.
@@
  • We propose a global routing flow using the proposed rule-based OARSMT algorithm to avoid obstacle violations at the early stage. This is followed by two obstacle-aware rip-and-reroute approaches, ultimately generating an obstacle-free solution for every net.
  • We propose a rule-based OARSMT algorithm that utilizes multiple rule-based schemes to enhance the results from FULTE and operates very quickly. With its high scalability, it can address most obstacle violations for all nets in a complete design during tree generation.
@@
  • The proposed obstacle-avoiding global routing eliminates obstacle violations while delivering a 1.96% improvement in wirelength and a 28.06% reduction in overflow cost on ISPD24 benchmarks with various obstacles, with limited via count and runtime overhead. Furthermore, The proposed OARSMT algorithm achieves a significant $\sim$10x-2700x and $\sim$150x-5800x runtime speedup compared to advanced OARSMT algorithms on randomized testcases and standard OARSMT benchmarks, respectively, with minimal wirelength overhead.
The remaining of this article is organized as follows: In Section II, a brief background and formulation on global routing and OARSMT is presented. In Section III, a brief introduction to our obstacle-avoiding global routing flow is proposed. In Section IV, the exact OARSMT algorithm in detail is described. In Section V, the details of our obstacle-avoiding global routing, including initial routing and rip-up and reroute, are presented. In Section VI, the experimental evaluation is discussed. Finally, in Section VII, the main conclusion is drawn.
@@ 2.1 Related Work
OARSMT Generation: OARSMT problem has been researched for many years. In the last century, the work [2] proposed the approach to calculate the shortest connection between two pins in a layout with obstacles. Recently, with the increasing constraints in new technologies, OARSMT becomes more important. For example, work [3] first constructs an RSMT without considering obstacles, and then uses a four-step algorithm to legalize the edges that intersect with the obstacles. However, the scenarios of the legalization is relatively straightforward and does not globally consider enough conditions. Work [4] proposes a solution using the obstacle-avoiding spanning graph (OASG), which trims edges and nodes to yield a minimum spanning tree for an acceptable quality OARSMT in polynomial time. However, it may overlook optimal solutions early on, resulting in excessive wirelength, and has impractical runtimes for complex cases. In contrast, work [5] also utilizes OASG but incorporates FLUTE for fine-tuning, achieving better quality and runtime, however it misses early global guidance from FLUTE. Work [6] approaches the OARSMT problem using the $\lambda$-geometry plane. Subsequent works [7, 8, 9, 10, 11] address multi-layer OARSMT, with work [11] being high quality by leveraging intermediate information of maze routing to optimize tree costs, reducing runtime while improving quality. Additionally, work [12] explores machine learning, employing a reinforcement learning algorithm to generate tree structures for small cases (12 pins, 20 obstacles) within limited runtimes. Compared with the previous works, leveraging the predefined rules, the proposed OARSMT algorithm has overall better quality, and has limited runtime, especially in large cases.
@@
To address the aforementioned issues, we propose a novel obstacle-avoiding global routing flow. This flow involves using our proposed fast and efficient OARSMT algorithm during the initial routing phase, and employing OARSMT-guided sparse maze routing and obstacle-aware sparse maze routing during the rip-up and reroute phase. Experiments indicate that this flow is effective when obstacles are widely involved.
@@ 2.2 Problem Formulation
In 3D global routing, the layout area is typically partitioned into an array of rectangular cells called global routing cells (GCells), which is represented as a 3D grid graph $G(V,E)$. Each GCell is treated as a vertex ($v\enspace\epsilon\enspace V$). A complete input case $C$ consists of lots of nets, with each net containing $m$ pins $\{p_{0},p_{1},p_{2}...,p_{m-1}\}$. Besides, there are $n$ obstacles $\{b_{0},b_{1},b_{2}...,b_{n-1}\}$ that are all rectangles with width $w_{i}$ and height $h_{i}$ for each $b_{i}$. The OARSMT problem and the global routing problem based on the obstacle-avoiding feature are defined as follows.
The OARSMT problem used in global routing tree generation is to find a rectilinear steiner tree with minimal wirelength that connects all pins, with no edge crossing any obstacles and no steiner node located inside any obstacles. The pins and obstacles are located in a $xy$-plane without overlap. Therefore, when executing OARSMT, it is necessary to map the obstacles in 3D space to 2D space. The cost of the tree is the wirelength, and an edge passing through (or a steiner node located at) the boundary of an obstacle is assumed to be legal.
For the complete global routing problem with obstacles, the capacity of an edge between a GCell represents the number of available routing tracks through it. The demand of an edge represents the capacity that has already been used on that specific edge. In the event where the demand of an edge exceeds its capacity, it leads to an overflow. It is established that obstacles are not traversable by route of net. As a consequence, the capacity of the area where obstacles are positioned is rendered as zero. The primary objective of obstacle-avoiding global routing is to minimize the $cost(C)=WL+\alpha_{ow}OW+\alpha_{ov}OV$, where $WL$ is the total wirelength of all nets, $OW$ is the total overflow cost, and $OV$ is the cost of obstacle violations. Note that since an obstacle is not traversable, its violation weight $\alpha_{ov}$ is high enough.
@@ 3 Proposed Global Routing Flow Overview
(1) OARSMT Generation: To minimize the number of net connections obstructed by obstacles during the initial routing, we construct an OARSMT in the tree generation stage. The OARSMT algorithm begins by using the FLUTE algorithm to determine the structure of the Rectilinear Steiner Minimum Tree (RSMT). Then, the algorithm updates the steiner nodes and edges that intersect with obstacles using a very fast newly proposed rule-based method, and completing the OARSMT generation. Then an initial route of the net is obtained in pattern routing. Compared to traditional initial routing methods, the proposed approach demonstrates a significant reduction in the number of nets that encounter obstacles, all while maintaining high processing speeds. This improvement can improve the subsequent obstacle-avoiding processes and offer valuable guidance for the later stages of routing.
@@
(3) Obstacle-aware Sparse Maze Routing: If there are still nets that have not successfully avoided obstacles, we will perform more refined maze routing for these nets called obstacle-aware sparse maze routing in this study. The obstacle-aware sparse grid graph will be constructed first, incorporating the boundaries of all obstacles, and maze routing will be performed within this graph. After this phase, the vast majority of nets can successfully navigate around obstacles.
@@ 4 OARSMT Algorithm Design
The first key process is the OARSMT generation during the initial routing. With great amount of nets, the proposed OARSMT algorithm can quickly generate the tree structures that avoid the obstacles. The generated OARSMTs can reduce the obstacle violations at the early step, which can release the workload of later steps and provide guidance for maze routing for better quality and runtime. In the following subsections, the design rationale of the proposed OARSMT algorithm is introduced, then the detailed algorithm and flow are elaborated.
@@ 4.1 Design Rationale and Edge Updating
As FLUTE does not consider obstacles, previous OARSMT works do not apply it in the early stage. This may leave out the valuable guidance of optimized tree structures, and incur more runtime. The proposed algorithm in this work leverages the solutions of FLUTE as a guidance, and further updates the solutions to optimize and legalize the layout with obstacles.
@@
The basic procedure to perform the solution updating is to find all the straight-line edges from a FLUTE solution that cross obstacles, and update them one by one. Using maze routing or recording the patterns of obstacles in a look-up table might be two approaches, but they either suffer from large runtime or tremendous patterns to be recorded. A trade-off approach can be the rule-based edge updating. Two example rules to update the edges from FLUTE are shown in Figure 2(a), including directly detouring the obstacles, and connecting through the obstacle boundaries. However, they can incur large cost or even violations with other obstacles, indicated as red lines in Figure 2(a).
@@
Figure 2: (a) Examples of directly detouring and connecting through the obstacle boundaries; (b) The examples where the proposed reference line-based edge updating show the optimal solutions; (c) An example where the steiner nodes generated by FLUTE are in an obstacle.
@@
To tackle the above challenges while keeping the fast process, a new rule-based edge updating approach is proposed, which keeps a balance between the two rules in Figures 2(a).
@@
In short, the new algorithm iteratively bends the edge line only when it is blocked on its direction, and each time bends to the obstacle corner nearest to the original edge. Taking the left part of Figures 2(b) as an example, the line of the original edge generated by FLUTE is named the reference line (such as $[n_{0},n_{1}]$). One of the nodes of the original edge is first selected as the source node (such as $n_{0}$), while another is the target node (such as $n_{1}$). The direction from the source node to the target node is named the reference direction, which is vertically from the top to the bottom in this example. Then, a line is created and extended from the source node to the target node through the reference direction. Once the line is blocked by an obstacle’s boundary (e.g., the line extended from $n_{0}$ is blocked at $n_{2}$ by the top boundary of $b_{0}$), it is bent and extended to a corner of this boundary (e.g., the extended line segment after bending is $[n_{2},c_{1}]$). The direction of bending is to the corner that is closer to the reference line (e.g., corner $c_{1}$ in this case). Once the bent line is extended and reaches the obstacle’s corner, it is then bent again and extends through the reference direction again (such as $[c_{1},n_{3}]$). Once the line is extended to have the same coordinate as the target node (e.g., $n_{3}$ has the same y-coordinate as $n_{1}$), the line is bent again towards the target node and connects it.
@@
Except the edge itself, the steiner nodes from FLUTE on edges also need to be updated when they are located inside the obstacles, such as the example in Figure 2(c). This can be fixed by regenerating the edges shown as the black lines in the figure, and removing the invalid edges and nodes. The black lines are generated as the shortest edges placed around the obstacle that can connect all the intersection points ($n_{i0}$-$n_{i2}$).

### v3

@@ Abstract
Routing is a crucial step in the VLSI design flow. With advancements in manufacturing technology, more constraints have emerged in design rules, particularly regarding obstacles during routing, leading to increased routing complexity. Unfortunately, many global routers struggle to generate efficient obstacle-free solutions due to the lack of scalable obstacle-avoiding tree generation methods and the capability to handle modern designs with complex obstacles and nets. In this work, we propose an efficient obstacle-aware global routing flow for VLSI designs with obstacles. The flow includes a rule-based obstacle-avoiding rectilinear Steiner minimal tree (OARSMT) algorithm during the tree generation phase. This algorithm is both scalable and fast, providing tree topologies avoiding obstacles in the early stage globally. With its guidance, in the later stages, the OARSMT-guided and obstacle-aware sparse maze routing are proposed to further minimize obstacle violations and reduce overflow costs. Compared to previously advanced methods on the benchmark with obstacles, our approach successfully eliminates obstacle violations and reduces wirelength and overflow cost, while sacrificing only a limited number of via counts and runtime overhead.
@@ 1 Introduction
As the number of cells and nets in modern designs has increased dramatically in advanced technology nodes, more complicated obstacles have emerged in global routing to accommodate more complex functionality requirements or to maintain consistency with earlier design phases. These obstacles may arise from various sources, including macro cells, intellectual property (IP) blocks, 3D packaging, pre-routed nets, power networks, etc. However, many global routers struggle to efficiently handle obstacle information and generate obstacle-free solutions at the global routing phase. High-efficiency obstacle-avoiding global routing has become more challenging and more necessary than ever before.
Global routing typically involves several steps, including tree generation, initial routing, and rip-up and reroute phases. Traditional routers that lack specialized methods for avoiding obstacles typically generate rectilinear Steiner minimal trees (RSMTs) for nets, which can be quickly obtained using established algorithms like FLUTE [1]. Without full consideration of obstacles, numerous design-rule violations can occur in the solutions. Although these violations can be addressed in later stages, with more time-consuming algorithms like maze routing, the overall efficiency and effectiveness of the routing process can become limited. To address the obstacle-avoiding RSMT (OARSMT) problem, recent research has introduced several algorithms for single-net solutions, such as complex graph algorithms, machine learning, etc. With more complex obstacle distributions, the complexity can significantly increase, especially for high fan-out nets. Additionally, the number of nets has also risen dramatically in modern designs. These factors result in scalability problems for most OARSMT algorithms in practicality when applied to large-scale cases in global routing. As a result, most previous research primarily concentrate on the tree generation phase and lack a comprehensive obstacle-avoiding global routing flow.
@@
  • We propose a global routing flow using the proposed rule-based OARSMT algorithm to avoid obstacle violations at the early stage. This is followed by two obstacle-aware rip-up and reroute approaches that ultimately generate obstacle-free solutions for all nets.
  • We propose a rule-based OARSMT algorithm that utilizes multiple rule-based schemes to enhance the results from FLUTE and operates very quickly. With its high scalability, it can address most obstacle violations for all nets in a complete design during tree generation.
@@
  • The proposed obstacle-avoiding global routing eliminates obstacle violations while achieving a 1.96% improvement in wirelength and a 28.06% reduction in overflow cost on ISPD24 benchmarks with various obstacles, with limited via count and runtime overhead, demonstrating high efficiency. Furthermore, the proposed OARSMT algorithm achieves significant speedups of $\sim$10x-2700x and $\sim$150x-5800x runtime compared to previously advanced OARSMT algorithms on randomized testcases and standard OARSMT benchmarks, respectively, with minimal wirelength overhead.
The remaining of this article is organized as follows: In Section II, a brief background and formulation on global routing and OARSMT is presented. In Section III, a brief introduction to our obstacle-avoiding global routing flow is proposed. In Section IV, the detailed OARSMT algorithm is described. In Section V, the details of our obstacle-avoiding global routing, including initial routing and rip-up and reroute, are presented. In Section VI, the experimental evaluation is discussed. Finally, in Section VII, the main conclusion is drawn.
@@ 2.1 Related Work
OARSMT Generation: The OARSMT problem has been researched for many years. In the last century, the work [2] proposed the approach to calculate the shortest connection between two pins in a layout with obstacles. Recently, with the increasing constraints in new technologies, OARSMT has become more important. For example, work [3] first constructs an RSMT without considering obstacles, and then uses a four-step algorithm to legalize the edges that intersect with the obstacles. However, the scenarios of the legalization are relatively straightforward and do not globally consider enough conditions. Work [4] proposes a solution using the obstacle-avoiding spanning graph (OASG), which trims edges and nodes to yield a minimum spanning tree for an acceptable quality OARSMT in polynomial time. However, it may overlook optimal solutions early on, resulting in excessive wirelength, and has impractical runtime for complex cases. In contrast, work [5] also utilizes OASG but incorporates FLUTE for fine-tuning, achieving better quality and runtime, but misses the early global guidance from FLUTE. Work [6] approaches the OARSMT problem using the $\lambda$-geometry plane. Subsequent works [7, 8, 9, 10, 11] address multi-layer OARSMT, with work [11] being high quality by leveraging intermediate information of maze routing to optimize tree costs, reducing runtime while improving quality. Additionally, work [12] explores machine learning, employing a reinforcement learning algorithm to generate tree structures for small cases (12 pins, 20 obstacles) within limited runtimes. Compared with the previous works, leveraging the predefined rules, the proposed OARSMT algorithm has overall better quality, and has limited runtime, especially in large cases.
@@
To address the aforementioned issues, we propose a novel obstacle-avoiding global routing flow. This flow involves using our proposed fast and efficient OARSMT algorithm during the initial routing phase, and employing OARSMT-guided sparse maze routing and obstacle-aware sparse maze routing during the rip-up and reroute phase. Experimental results demonstrate that this flow is effective when obstacles are widely involved.
@@ 2.2 Problem Formulation
In 3D global routing, the layout area is typically partitioned into an array of rectangular cells called global routing cells (GCells), which is represented as a 3D grid graph $G(V,E)$. Each GCell is treated as a vertex ($v\enspace\epsilon\enspace V$). A complete input case $C$ consists of lots of nets, with each net containing $m$ pins $\{p_{0},p_{1},p_{2},...,p_{m-1}\}$. Additionally, there are $n$ obstacles $\{b_{0},b_{1},b_{2},...,b_{n-1}\}$ that are all rectangles with width $w_{i}$ and height $h_{i}$ for each $b_{i}$. The OARSMT problem and the global routing problem based on the obstacle-avoiding feature are defined as follows.
Following the typical previous works [4, 11, 21], the OARSMT problem used in global routing tree generation is to find a rectilinear Steiner tree with minimal wirelength that connects all pins, with no edge crossing any obstacles and no node located inside any obstacles. Note that edges passing through the boundaries are legal. The pins and obstacles are located on a $xy$-plane without overlap. Therefore, when executing OARSMT, it is necessary to map the obstacles in 3D space to 2D space. The cost of the tree is the wirelength, and an edge passing through (or a node located at) the boundary of an obstacle is assumed to be legal.
For the complete global routing problem with obstacles, the capacity of an edge between two adjacent GCells represents the number of available routing tracks through it. The demand of an edge represents the capacity that has already been used on that specific edge. If the demand of an edge exceeds its capacity, it results in an overflow. It is established that obstacles are not traversable by route of net. As a consequence, the capacity of the area where obstacles are positioned is set to zero. The primary objective of obstacle-avoiding global routing is to minimize the $cost(C)=WL+\alpha_{ow}OW+\alpha_{ov}OV$, where $WL$ is the total wirelength of all nets, $OW$ is the total overflow cost, and $OV$ is the cost of obstacle violations. Note that since an obstacle is not traversable, its violation weight $\alpha_{ov}$ is sufficiently high.
For a better description, we use $(x,y)$ to represent a coordinate. Meanwhile, $n_{i},n_{j}]$ represents a straight line between node $i$ and $j$. $n_{i}.x$ and $n_{i}.y$ stand for the x and y coordinates of node $i$. Moreover, $n_{i},dir)$ represents a ray segment line from node $n_{i}$ with direction dirdir (A ray segment is a straight line extending from a point indefinitely in one direction only).
@@ 3 Proposed Global Routing Flow Overview
(1) OARSMT Generation: To minimize the number of net connections obstructed by obstacles during the initial routing, we construct an OARSMT in the tree generation stage. The OARSMT algorithm begins by using the FLUTE algorithm to determine the structure of the Rectilinear Steiner Minimum Tree (RSMT). Then, the algorithm updates the nodes (steiner nodes and L-shape corner nodes) and edges that intersect with obstacles using a very fast newly proposed rule-based method, and completes the OARSMT generation. Then, an initial route of the net is obtained through pattern routing. Compared to traditional initial routing methods, the proposed approach demonstrates a significant reduction in the number of nets that encounter obstacles, all while maintaining high processing speeds. This improvement can enhance the subsequent obstacle-avoiding processes and offer valuable guidance for the later stages of routing.
@@
(3) Obstacle-aware Sparse Maze Routing: If there are still nets that have not successfully avoided obstacles, we will perform more refined maze routing for these nets, referred to as obstacle-aware sparse maze routing in this study. The obstacle-aware sparse grid graph will be constructed first, incorporating the boundaries of all obstacles, and maze routing will be performed within this graph. After this phase, the vast majority of nets are able to successfully avoid obstacles.
@@ 4 OARSMT Algorithm Design
The first key process is the OARSMT generation during the initial routing. With a large number of nets, the proposed OARSMT algorithm can quickly generate tree structures that avoid obstacles. The generated OARSMTs can reduce the obstacle violations at the early step, which can release the workload of later steps and provide guidance for maze routing to achieve better quality and runtime. In the following subsections, the design rationale of the proposed OARSMT algorithm is introduced, followed by a detailed elaboration of the algorithm and flow.
@@ 4.1 Design Rationale and Edge Updating
As FLUTE does not consider obstacles, previous OARSMT works do not apply it in the early stage. This may leave out the valuable guidance of optimized tree structures, and incur more runtime. The proposed algorithm in this work leverages the solutions of FLUTE as guidance, and further updates them to optimize and legalize the layout with obstacles.
@@
The basic procedure to perform the solution updating is to find all the straight-line edges based on a FLUTE solution (with L-shapes in both directions considered for two nodes that a single horizontal or vertical line cannot connect) that cross obstacles, and update them one by one. Using maze routing or recording the patterns of obstacles in a look-up table might be two approaches, but they either suffer from long runtime or tremendous patterns to be recorded. A trade-off approach can be the rule-based edge updating. Examples about two straightforward rules to update the edges are shown in Figure 2(a)-(c), including directly detouring around the obstacles, and connecting through the obstacle boundaries. However, directly detouring is too aggressive, while connecting through obstacle boundaries is too conservative. They can incur violations or large cost, indicated as red lines in Figure 2(b)-(c).
@@
Figure 2: (a) An example of directly detouring; (b) An example where directly detouring is invalid, while connecting through obstacle boundaries is optimal; (c) An example where connecting through obstacle boundaries incurs large cost; (d) The examples where the proposed reference line-based edge updating show the optimal solutions; (e) An example where both the Steiner node and L-shape corner node are located inside an obstacle.
Algorithm 1 Edge updating algorithm
Function: Edge_Update
Input: Source, Target Nodes $n_{s}$, $n_{t}$; Reference Line $l_{r}$; Obstacles $S_{b}$.
Output: Updated Edges $U_{e}$.
[…]
4:while True do
5:  $b_{1}\leftarrow$ The 1st obstacle $\in S_{b}$ that blocks $n^{\to},rdir)$
6:  if $b_{1}!=NULL$ then
7:   $c_{n0},c_{n1}]\leftarrow$ $b_{1}$’s 1st boundary that blocks $n^{\to},rdir)$
[…]
9:   $U_{e}\leftarrow U_{e}\bigcup\{[n^{\to},n^{\dashv}]\}$
10:   $n^{\to}\leftarrow n^{\dashv}$
11:   $n^{\dashv}\leftarrow dis(c_{n0},l_{r})<dis(c_{n1},l_{r})?c_{n0}:c_{n1}$
12:   $U_{e}\leftarrow U_{e}\bigcup\{[n^{\to},n^{\dashv}]\}$
13:   $n^{\to}\leftarrow n^{\dashv}$
14:  else
[…]
16:   $U_{e}\leftarrow U_{e}\bigcup\{[n^{\to},n^{\dashv}]\}$
17:   $U_{e}\leftarrow U_{e}\bigcup\{[n^{\dashv},n_{t}]\}$
18:   Break
19:  end if
20:end while
21:Return $U_{e}$
@@
To tackle the above challenges while keeping the fast process, a new rule-based edge updating approach is proposed in Algorithm 1, which keeps a balance between the two rules in Figure 2(a)-(c).
@@
We first show an example in Figure 2(d). To update an edge $n_{0},n_{1}]$ that crosses obstacles, the proposed algorithm extends the edge line from $n_{0}$ to $n_{1}$, and iteratively bends the line when it is blocked on its extending direction. The edge always bends toward the obstacle corner nearest to the original edge. Taking the left part of Figure 2(d) as an example, the line of the original edge $n_{0},n_{1}]$ based on FLUTE is named the reference line. The nodes of the original edge $n_{0},n_{1}]$ are named source node $n_{0}$ and target node $n_{1}$, respectively. The direction from the source node to the target node is named the reference direction, which is vertical down in this example. A line is initially created and extended from $n_{0}$ to $n_{1}$ along the reference direction. Once the extended line ($n_{0},n_{2}]$) is blocked by an obstacle’s boundary ($c_{0},c_{1}]$ of $b_{0}$), it is bent and extended to the boundary’s corner ($c_{1}$) that is closer to the reference line, resulting in a line segment $n_{2},c_{1}]$. Then, it is bent again to extend through the reference direction, resulting in $c_{1},n_{3}]$. Once the line is extended to reach the same coordinate on the reference direction as the target node $n_{1}$ (e.g., $n_{3}$ has the same y-coordinate as $n_{1}$), the line is bent again towards $n_{1}$ and connects it. This edge from the final bending (such as $n_{3},n_{1}]$) might still cross a few obstacles, which can be solved by iteratively performing the edge updating.
The detailed algorithm is shown as Algorithm 1, which can update edge $n_{s},n_{t}]$ according to the reference line $l_{r}$, given obstacles $S_{b}$, and output the updated edges $U_{e}$ (For better description, the information of the nodes is also assumed included in the information of edges). During initialization, $U_{e}$ is set to be empty. A node $n^{\to}$ is maintained to be the current node to be extended from. It is initialized as $n_{s}$ at Line 2 of Algorithm 1. The reference direction rdirrdir is initialized at Line 3. Next, similar to the example in Figure 2(d), the edge is extended through rdirrdir like a ray segment $n^{\to},rdir)$ until blocked by an obstacle $b_{1}$ (Line 5). Then, the first boundary of $b_{1}$ that blocks $n^{\to},rdir)$ is extracted (Line 7). For example, in Figure 2(d), when $n^{\to}$ is $n_{0}$, $c_{n0}$$c_{n1}]$ is $c_{0}$$c_{1}]$. Then, $n^{\dashv}$ is the node where the current edge extension is end. If the reference direction is horizontal, $n^{\dashv}$’s coordinate contains $c_{n0}$’s x-coordinate and $n^{\to}$’s y-coordinate. Otherwise, $n^{\dashv}$’s coordinate contains $n^{\to}$’s x-coordinate and $c_{n0}$’s y-coordinate (Line 8). Now, current edge extension result $n^{\to},n^{\dashv}]$ is recorded into $U_{e}$ (Line 9). The next part is to bend the edge extension to a $b_{1}$’s corner that is closer to $l_{r}$, and record the bent line segment as a new edge (Lines 10-12). Finally, $n^{\to}$ is updated to the reached corner and the procedure is performed again (Line 13). When the edge extension from $n^{\to}$ is never blocked by further obstacles, the edge is extended to the $n^{\dashv}$ node with the same coordinate in the reference direction as $n_{t}$ (Line 15). The current edge is then recorded (Line 16), and the edge from the $n^{\dashv}$ to the target node is finally recorded to finish the connection (Line 17). The edge updating is finished (Line 18).
@@
In addition to the edge itself, both the Steiner nodes and the L-shape corner nodes along the edges also need to be updated when they are located inside the obstacles, such as the example in Figure 2(e). This can be fixed by regenerating the edges shown as the black lines in the figure, and removing the invalid edges and nodes. The black lines are generated as the shortest edges placed around the obstacle that can connect all the intersection points ($n_{i0}$-$n_{i2}$).